# These $\altaffiliationThese authors contributed equally to this work.$

### v1

@@ 5 Basis-Set Generation Recipe
$(ia|P)_{m}=(P|ia)_{m}=\int d\mathbf{r}\;d\mathbf{r}^{\prime}\;\phi_{i}(\mathbf{r})\;\phi_{j}(\mathbf{r})\;m(\mathbf{r},\mathbf{r}^{\prime})\;\varphi_{P}(\mathbf{r}^{\prime})\,$,
@@
Here, we introduced the auxiliary RI basis set $\{\varphi_{P}\}$, which also consists of Gaussians. $m(\mathbf{r},\mathbf{r}^{\prime})$ is the RI metric; convergence of the RI expansion (LABEL:e28) depends on $m$. It has been shown that the fastest convergence of the RI expansion is achieved using the Coulomb metric, $m(\mathbf{r},\mathbf{r}^{\prime})\hskip 1.69998pt{=}\hskip 1.69998pt1/|\mathbf{r}\hskip 1.69998pt{-}\hskip 1.69998pt\mathbf{r}^{\prime}|$58.
Early applications of RI include DFT 59, 60 and MP2 61, where it became a standard technique by now. In random phase approximation (RPA) calculations for the correlation energy, RI reduces the scaling from $O(N^{6})$ in the canonical Casida-based formulation to $O(N^{4})$ 62. However, RI is not universally beneficial: in Hartree–Fock and hybrid functional calculations, RI typically improves performance only when large orbital basis sets are used 63. For small orbital basis sets, conventional four-center formulations may remain more efficient. For the computation of charged excitations based on $GW$, RI has also become a common tool, where it reduces the scaling from $O(N^{6})$ to $O(N^{4})$ 64, 65, 66, as well as for charge-neutral excitations based on the BSE, where the screened Coulomb interaction is computed using RI.
Figure 3: $\Delta_{\mathrm{I}}$ metric as a function of the auxiliary RI basis set size, using various augmented MOLOPT basis sets introduced in this work. The auxiliary RI basis sets are optimized for a given basis set size to match the MP2 correlation energy. Reference auxiliary RI basis set sizes and $\Delta_{\mathrm{I}}$ for aug-cc-pVDZ-RIFIT and aug-$c$-pVTZ-RIFIT are shown for comparison.
When using RI, an auxiliary RI basis set $\{\varphi_{P}\}$ is required for the factorization (LABEL:e28) of four-center integrals into two- and three-center integrals. Although it is possible to generate auxiliary RI basis sets on the fly during the calculation 67, this often results in large auxiliary RI basis sets. Instead, we generate auxiliary RI basis sets by matching the RI-MP2 correlation energy of isolated atoms to the corresponding MP2 reference energies 68. For this purpose, as proposed in 68, we generate auxiliary RI basis sets of different sizes by using the (relative) $\Delta_{\mathrm{I}}$ metric as an optimization parameter:
@@
where the $(i,j)$ refer to occupied orbitals and $(a,b)$ to empty orbitals, $E_{\mathrm{MP2}}$ is the MP2 correlation energy and using a standard notation for the double bar integral defined as 69:
@@
For comparison, we compute $\Delta_{\mathrm{I}}$ for the existing aug-cc-pVDZ and aug-cc-pVTZ with corresponding RI basis sets 70, see Fig. 3. These basis sets have very small $\Delta_{\mathrm{I}}$ metric values below $10^{-6}$, but are relatively large in size (72 and 106 functions for aug-cc-pVDZ-RIFIT and aug-cc-pVTZ-RIFIT, respectively). We also create smaller auxiliary RI basis sets with lower accuracy, which are still sufficient in applications as we demonstrate later for nanographenes (Sec. 6). All generated auxiliary RI basis sets are available in the Supporting Information (Sec. S6). The optimization was performed using the auxiliary RI basis set optimizer implemented in CP2K 71.
7 Description of molecular test set
For benchmarking excited-state energies with our generated basis sets, we use the GWGW5000 dataset 72. We exclude all molecules with less than ten atoms as small molecules tend to have very diffuse unoccupied states; and the purpose of the generated basis sets is to describe large molecules and crystals with less diffuse unoccupied states. To reduce the computational cost, we only use molecules with at most 20 atoms. We also remove all molecules larger than 15 atoms in which carbon atoms outnumber all other non-hydrogen elements by more than a factor of two. Our aim is to ensure a balanced benchmark set avoiding overrepresentation of unsubstituted or weakly substituted hydrocarbons. The precise criterion for removal is $N_{\text{C}}\hskip 1.69998pt{>}\hskip 1.69998pt2(N_{\text{tot}}{-}N_{\text{H}}{-}N_{\text{C}}),$ where $N_{\text{C}}$ is the number of carbon atoms, $N_{\text{H}}$ the number of hydrogen atoms and $N_{\text{tot}}$ the total number of atoms in the molecule.
Applying these criteria gives 247 molecules in the GWGW5000 benchmark set, where the majority of the molecules contain C (98 %), H (96 %), N (76 %), O (74 %), while other elements are less often present: S (31 %), Cl (23 %), F (10 %), P (2 %), B (1 %) and Si (1 %). We also use a second molecular benchmark set that focuses on other elements (Li, Be, B, Na, Ca, Al, Si, P); we show the composition of this benchmark set and the calculations in the supplementary information SI1 and SI2, respectively.
8 Computational details
We employ the CP2K package for all calculations 41. CP2K employs a Gaussian basis set for representing KS orbitals [Eq. (2)]. We use the Gaussian and augmented plane-waves scheme 73, which enables all-electron calculations in CP2K. We use implementations in CP2K of conventional $GW$ (Sec. 9) in imaginary-frequency formulation with analytic continuation 45, BSE and TDDFT 74 (Sec. 10), as well as low-scaling $GW$ 75 (Sec. 1112) based on the space-time method 16 using minimax time-frequency grids 76, 77, 78, 79.
We employ the PBE0 exchange-correlation functional 80 as starting point for our excited-state calculations. The usage of PBE0 as starting point for $GW$ and Bethe-Salpeter avoids numerical instabilities due to multipole features of the self-energy close to the quasiparticle solution 81, 82, which can be present when starting from the PBE functional 83. For the low–scaling $GW$ calculations on nanographenes (Sec. 12), however, we use PBE for the SCF cycle to reduce computational cost. As discussed in Ref.82, $G_{0}W_{0}$@PBE can suffer from numerical instabilities caused by poles in the self-energy $\Sigma_{n}(\omega)$ close to the quasiparticle energy $\omega\hskip 1.69998pt{\approx}\hskip 1.69998pt\varepsilon_{n}^{G_{0}W_{0}}$, where
$\varepsilon_{n}^{G_{0}W_{0}}=\varepsilon_{n}^{\text{PBE}}+\text{Re},\Sigma_{n}(\varepsilon_{n}^{G_{0}W_{0}})-v_{n}^{\text{xc}}.$ (32)
@@ 8 Computational details
leading to the modified quasiparticle equation
@@
9 PBE0 and GW HOMO-LUMO gaps
We compute HOMO-LUMO gaps using PBE0 and $G_{0}W_{0}@$PBE0 for a subset of 247 molecules from the GWGW5000 benchmark set (see Sec. 7). Fig. 4a,b shows the results for the aug-MOLOPT basis sets introduced in this work, together with the aug-cc-pVXZ, cc-pVXZ, and MOLOPT basis sets.
where $N_{\text{mol}}\hskip 1.69998pt{=}\hskip 1.69998pt247$ is the number of molecules and $\Delta_{i}^{B}$ is the PBE0 HOMO-LUMO gap of molecule $i$ computed with basis set $B$. While MAD captures the average accuracy, we also report the 95th percentile error (95PE) to quantify the worst-case deviations of the worst 5% of molecules. The aug-MOLOPT basis sets show systematic improvement from aug-SZV-MOLOPT-ae to aug-TZVP-MOLOPT-ae, with both MAD and 95PE decreasing towards the CBS limit. The aug-TZVP-MOLOPT basis has a MAD of just 14 meV. The aug-cc-pVXZ basis sets generally show even smaller deviations at equivalent basis size—e.g., aug-cc-pVTZ is closer to the CBS than aug-TZVP-MOLOPT-ae. This is expected, as aug-cc-pVXZ are specifically optimized for electron affinities and include very diffuse functions well suited for describing the LUMO. In contrast, the cc-pVXZ and MOLOPT families exhibit significantly larger errors and slower convergence for the HOMO-LUMO gap, reflecting their optimization for ground-state energies rather than excited states. Overall, the aug-MOLOPT basis sets provide fast convergence of HOMO-LUMO gaps, while maintaining a well-conditioned overlap matrix (see Fig. 1).
Fig. 4b shows the basis set convergence of the four basis set families for $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps, using aug-cc-pV5Z as the CBS reference. The aug-MOLOPT basis sets exhibit consistently small deviations from the CBS. Notably, the MAD of the small aug-SZV-MOLOPT-ae basis is 160 meV, better than the larger aug-cc-pVDZ basis, which yields a MAD of 220 meV. Likewise, aug-DZVP-MOLOPT-ae achieves a MAD of just 60 meV, below the 80 meV deviation of the larger aug-cc-pVTZ basis. This finding suggests that the aug-MOLOPT basis sets are the ideal choice for $GW$ HOMO-LUMO gap calculations to ensure both fast basis set convergence of $GW$ HOMO-LUMO gaps and numerical stability for large molecules and crystals. Again, the non-augmented cc-pVXZ and MOLOPT basis sets exhibit larger and more slowly converging errors for $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps. For example, the MAD of the large cc-pV5Z basis is 240 meV—only slightly lower than the minimal aug-SZV-MOLOPT-ae-mini basis which has a MAD of 270 meV. When excluding molecules with diffuse LUMOs (defined as LUMO eigenvalues above – 2 eV), the MAD for cc-pV5Z decreases by an order of magnitude to 30 meV (Fig. S2 in supporting information). This indicates that basis sets optimized for ground-state properties, and lacking diffuse functions, are inadequate for accurately describing $GW$ gaps in systems with unoccupied states that significantly extend into the vacuum.
10 GW+BSE and TDDFT excitation energies
[…]
Figure 4d shows the basis set convergence of the first ten excitation energies computed with TDDFT (PBE0). As before, the aug-MOLOPT basis sets exhibit systematic improvement with increasing basis size. However, in contrast to the BSE case, the aug-cc-pVXZ basis sets outperform the aug-MOLOPT family: for example, the MAD of aug-DZVP-MOLOPT-ae is 100 meV, whereas aug-cc-pVDZ achieves a significantly lower MAD of 17 meV. Comparing with BSE results in Fig. 4c, the aug-cc-pVXZ basis sets converge more rapidly for TDDFT than for BSE, while the aug-MOLOPT sets show similarly fast convergence for both methods. We attribute this difference to the design philosophy behind the basis sets: the aug-MOLOPT sets were specifically optimized for BSE excitation energies (albeit on a different training set, the Thiel’s set 86), whereas the aug-cc-pVXZ family was not. Nevertheless, aug-cc-pVXZ basis sets feature ill-conditioned overlap matrices for large molecules, making the aug-MOLOPT basis sets numerically more robust for larger molecules and crystals.
[…]
For the $GW$ and BSE basis set benchmark presented in Fig. 4, we employed the largest available auxiliary RI basis sets (see Sec. 6 for generation details). To enable large-scale $GW$ and BSE simulations, it is desirable to reduce the size of the auxiliary RI basis set while maintaining high numerical accuracy. Smaller auxiliary RI basis sets lead to lower computational cost and improved scalability, particularly in low-scaling $GW$ algorithms.
In this work, we employ the cubic-scaling $GW$ implementation in CP2K for molecules 75, which uses the truncated Coulomb metric 87 for the RI approximation. While the fastest convergence of RI-based methods is achieved when the cutoff radius of the Coulomb operator is infinite, this also increases the computational cost. Therefore, a balance must be found: the cutoff radius should be small enough to reduce computational requirements, yet large enough to ensure sufficiently fast convergence of the auxiliary RI basis set.
To evaluate this tradeoff, we benchmark $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps for the aug-SZV-MOLOPT-ae basis set on the same subset of 247 molecules from the GWGW5000 benchmark set used in Fig. 4. We consider four auxiliary RI basis sets of increasing size, corresponding to decreasing the $\Delta_{\mathrm{I}}$ metric threshold: $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$. For each basis set, we evaluate four different cutoff values for the truncated Coulomb operator: $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt3$, 5, 7, and 9 Å.
Figure 5: Convergence of low-scaling $GW$ calculations 75 with respect to the cutoff radius of the truncated Coulomb metric and the auxiliary RI basis set size (here quantified by the $\Delta_{\mathrm{I}}$ metric threshold). As orbital basis set, we employ aug-SZV-MOLOPT-ae. The color map shows the mean absolute deviation of $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps for the same subset of 247 molecules from the GWGW5000 benchmark set used in Fig. 4, relative to a reference calculation using the aug-cc-pV5Z-RIFIT auxiliary RI basis set $8$. Each row corresponds to an auxiliary RI basis set generated with a given $\Delta_{\mathrm{I}}$ metric threshold (from $10^{-2}$ to $10^{-5}$). Smaller errors are achieved with tighter RI thresholds and larger Coulomb cutoffs. A practical compromise is reached with a $\Delta_{\mathrm{I}}$ metric threshold of $10^{-4}$ and $r_{c}\hskip 1.69998pt{\geq}\hskip 1.69998pt7$ Å (error: 30 meV).
Fig. 5 summarizes the results. The color map shows the absolute deviation of the $G_{0}W_{0}$ HOMO-LUMO gaps (averaged over all 247 molecules) from the converged reference obtained with the large aug-cc-pV5Z-RIFIT auxiliary RI basis set 88 and cutoff $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt9$ Å. At the loosest RI threshold ($10^{-2}$) and smallest cutoff ($r_{c}=3$ Å), the average error exceeds 300 meV. However, increasing the cutoff to $r_{c}=9$ Å reduces this error to $\sim$ 130 meV. For a larger auxiliary RI basis ($\Delta_{\mathrm{I}}$ threshold of $10^{-4}$), numerical accuracy is substantially improved: for $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt7$ Å, the deviation is 30 meV, and drops to 20 meV at $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt9$ Å. The best overall agreement with aug-cc-pV5Z-RIFIT is obtained for an even larger auxiliary RI basis ($\Delta_{\mathrm{I}}$ threshold of $10^{-5}$) with $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt9$ Å, where the average absolute error is reduced to 13 meV.
These results demonstrate that accurate low-scaling $GW$ calculations can be achieved using relatively compact auxiliary RI basis sets when paired with a sufficiently large Coulomb cutoff. For practical applications aiming at high numerical precision, we recommend a $\Delta_{\mathrm{I}}$ threshold of $10^{-4}$ and a cutoff radius of at least 7 Å, giving excellent balance between efficiency and accuracy ($\sim$ 30 meV).
Figure 6: GWGW calculations on nanographenes of increasing length (defined by the number of stripes $L\hskip 1.69998pt{=}\hskip 1.69998pt1,2,4,\dots,128$). (a) Nanographene geometry for $L\hskip 1.69998pt{=}\hskip 1.69998pt1,2,4,8$. Note that we put two hydrogen atoms at the center carbon atom at the zigzag edge to prevent for magnetic zigzag edge states. (b) Quasiparticle HOMO-LUMO gaps computed with $G_{0}W_{0}$ + Hedin’s shift @ PBE using different basis sets. Inset: deviation from the aug-DZVP-MOLOPT-ae basis set. (c) PBE LUMO eigenvalue, serving as a measure of LUMO diffuseness. (d) Condition number $\kappa(\mathbf{S})$ of the overlap matrix, computed from Eq. (11). (e) Computation time (in core hours) of the $G_{0}W_{0}$ calculations on Noctua2 (AMD Milan 7763). The aug-MOLOPT basis sets exhibit low condition numbers and reduced computational cost, enabling stable and efficient calculations for nanographenes exceeding 2000 atoms. Details on the number of nodes used, wall time and memory consumption of the $GW$ calculations are listed in Table 3.
[…]
@@ 10 GW+BSE and TDDFT excitation energies
To demonstrate the suitability of the generated augmented MOLOPT basis sets for large-scale applications, we perform $GW$ calculations on nanographenes of increasing size. Representative geometries are shown in Fig. 6a. For these systems, we employ the PBE functional 83 as the DFT starting point and Hedin’s shift (34) to avoid the higher cost of hybrid functionals during the SCF.
@@
We use auxiliary RI basis sets with a $\Delta_{\mathrm{I}}$ value below $1.5\cdot 10^{-2}$. The corresponding basis sizes are listed in Table 2. $G_{0}W_{0}$+H HOMO-LUMO gaps for the nanographenes are shown in Fig. 6b. For $L\hskip 1.69998pt{=}\hskip 1.69998pt1$ (9,10-Dihydroanthracene), basis set convergence is challenging: the minimal aug-SZV-MOLOPT-ae-mini basis underestimates the gap by approximately 1 eV. This is partly due to the small auxiliary RI basis used (cf. Fig. 5). For larger systems ($L\geq 8$), convergence improves significantly: all five basis sets agree within 50 meV (inset of Fig. 6b). This matches findings for twodimensional materials 23, where convergence within 100 meV was reached using the aug-SZV-MOLOPT basis.
We attribute the improved basis set convergence for larger structures to three factors: (i) larger systems offer more basis functions, increasing flexibility; (ii) the PBE LUMO energy decreases with $L$ (Fig. 6c), making the LUMO less diffuse in vacuum and thus easier to represent (as discussed for the decay length $\zeta_{n}\hskip 1.69998pt{\approx}\hskip 1.69998pt\hbar/{\sqrt{2m|\varepsilon_{n}|}}$ in Sec. 5222For a benchmark of the numerical precision of basis sets for $G_{0}W_{0}$ HOMO-LUMO gaps as function of the DFT LUMO energy, see the supporting information, Fig. S1 and S2 (and Fig. S3 for differences of excitation energies).; (iii) there may be cancellation of errors between an underconverged orbital and auxiliary RI basis set. Notably, error cancellation does not distort the size dependence: for $L\hskip 1.69998pt{=}\hskip 1.69998pt32$, 64, 128, all aug-SZV-MOLOPT (SR, mini) basis sets yield size-converged gaps consistently within 33 meV. This indicates robust $GW$ calculations for large systems. These results support two conclusions: (i) basis set convergence for nanostructures differs from that of small molecules and must be analyzed accordingly; (ii) further optimization of Gaussian basis sets for extended systems, in particular with pseudopotentials 89, appear promising.
The condition number of the overlap matrix remains below $10^{7}$ for all augmented MOLOPT basis sets (Fig. 6d). In contrast, it exceeds $10^{13}$ for aug-cc-pVDZ. The computational cost is roughly reduced by a factor of 280 when using aug-SZV-MOLOPT-ae-mini instead of aug-cc-pVDZ (Fig. 6e). This aligns with expected $GW$ scaling of $N_{\text{bf}}^{2}N_{\text{RI}}^{2}$: According to Table 2, $N_{\text{bf}}$ and $N_{\text{RI}}$ decrease by factors of about 2.4 and 6.8, respectively, giving $2.4^{2}\hskip 1.19995pt{\cdot}\hskip 1.19995pt6.8^{2}\hskip 1.69998pt{\approx}\hskip 1.69998pt270$. Despite this enormous speedup, the gap difference between aug-SZV-MOLOPT-ae-mini and aug-cc-pVDZ is less than 10 meV for $L\hskip 1.69998pt{=}\hskip 1.69998pt16$.
[…]

### v2

@@ 3.1 Auxiliary RI basis set generation
$(ia|P)_{m}=(P|ia)_{m}=\int d\mathbf{r}\;d\mathbf{r}^{\prime}\;\psi_{i}(\mathbf{r})\;\psi_{a}(\mathbf{r})\;m(\mathbf{r},\mathbf{r}^{\prime})\;\varphi_{P}(\mathbf{r}^{\prime})\,$,
@@
Here, we introduced the auxiliary RI basis set $\{\varphi_{P}\}$, which also consists of Gaussians. $m(\mathbf{r},\mathbf{r}^{\prime})$ is the RI metric; convergence of the RI expansion (LABEL:e28) depends on $m$. It has been shown that the fastest convergence of the RI expansion is achieved using the Coulomb metric, $m(\mathbf{r},\mathbf{r}^{\prime})\hskip 1.69998pt{=}\hskip 1.69998pt1/|\mathbf{r}\hskip 1.69998pt{-}\hskip 1.69998pt\mathbf{r}^{\prime}|$ Vahtras1993.
Early applications of RI include DFT Eichkorn1995, Sierka2003 and MP2 weigend2003, where it became a standard technique by now. In random phase approximation (RPA) calculations for the correlation energy, RI reduces the scaling from $O(N^{6})$ in the canonical Casida-based formulation to $O(N^{4})$ Eshuis2010. However, RI is not universally beneficial: in Hartree–Fock and hybrid functional calculations, RI typically improves performance only when large orbital basis sets are used Weigend2002. For small orbital basis sets, conventional four-center formulations may remain more efficient. For the computation of charged excitations based on $GW$, RI has also become a common tool, where it reduces the scaling from $O(N^{6})$ to $O(N^{4})$ blase2011, Ren2012, van2013gw, as well as for charge-neutral excitations based on the BSE, where the screened Coulomb interaction is computed using RI.
Figure 4: $\Delta_{\mathrm{I}}$ metric for a carbon atom as a function of the auxiliary RI basis set size, using various augmented MOLOPT basis sets introduced in this work. The auxiliary RI basis sets are optimized for the carbon atom in a given basis set size to match the MP2 correlation energy. Reference auxiliary RI basis set sizes and $\Delta_{\mathrm{I}}$ for aug-cc-pVDZ-RIFIT and aug-$c$-pVTZ-RIFIT are shown for comparison.
When using RI, an auxiliary RI basis set $\{\varphi_{P}\}$ is required for the factorization (LABEL:e28) of four-center integrals into two- and three-center integrals. Although it is possible to generate auxiliary RI basis sets on the fly during the calculation Stoychev2017, Lehtola2021, this often results in large auxiliary RI basis sets. Recently, several schemes have been proposed to tackle this issue, such as the combined use of a contraction based on the singular value decomposition and a high-momentum truncation of the generated basis sets Lehtola2023, or a newly suggested approach with uncontracted, non-even-tempered sets that are truncated using the 2-body energy as a metric Valeev2025. In this work, we instead generate auxiliary RI basis sets by matching the RI-MP2 correlation energy of isolated atoms to the corresponding MP2 reference energies Weigend1998. For this purpose, as proposed in Weigend1998, we generate auxiliary RI basis sets of different sizes by using the (relative) $\Delta_{\mathrm{I}}$ metric as an optimization parameter:
@@
where the $(i,j)$ refer to occupied orbitals and $(a,b)$ to empty orbitals, $E_{\mathrm{MP2}}$ is the MP2 correlation energy and using a standard notation for the double bar integral defined as Kraka1991:
@@
For comparison, we compute $\Delta_{\mathrm{I}}$ for the existing aug-cc-pVDZ and aug-cc-pVTZ with corresponding RI basis sets weigend2002a, see Fig. 4. These basis sets have very small $\Delta_{\mathrm{I}}$ metric values below $10^{-6}$, but are relatively large in size (72 and 106 functions for aug-cc-pVDZ-RIFIT and aug-cc-pVTZ-RIFIT, respectively). We also create smaller auxiliary RI basis sets with lower accuracy, which are still sufficient in applications as we demonstrate later for nanographenes (Sec. 6). All generated auxiliary RI basis sets are available in the Supporting Information (Sec. S6). The optimization was performed using the auxiliary RI basis set optimizer implemented in CP2K DelBen2013.
@@
For the $GW$ and BSE basis set benchmark presented in Fig. 3, we employed the largest available auxiliary RI basis sets (see Sec. 3.1 for generation details). To enable large-scale $GW$ and BSE simulations, it is desirable to reduce the size of the auxiliary RI basis set while maintaining high numerical accuracy. Smaller auxiliary RI basis sets lead to lower computational cost and improved scalability, particularly in low-scaling $GW$ algorithms.
In this work, we employ the cubic-scaling $GW$ implementation in CP2K for molecules Wilhelm2021, which uses the truncated Coulomb metric Jung2005 for the RI approximation. While the fastest convergence of RI-based methods is achieved when the cutoff radius of the Coulomb operator is infinite, this also increases the computational cost. Therefore, a balance must be found: the cutoff radius should be small enough to reduce computational requirements, yet large enough to ensure sufficiently fast convergence of the auxiliary RI basis set.
To evaluate this trad-eoff, we benchmark $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps for the aug-SZV-MOLOPT-ae, aug-DZVP-MOLOPT-ae and aug-TZVP-MOLOPT-ae basis sets on the same subset of 247 molecules from the GWGW5000 benchmark set used in Fig. 3. We consider four auxiliary RI basis sets of increasing size, corresponding to decreasing the $\Delta_{\mathrm{I}}$ metric threshold: $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$. For each basis set, we evaluate four different cutoff values for the truncated Coulomb operator: $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt3$, 5, 7, and 9 Å.
Figure 5: Convergence of low-scaling $GW$ calculations Wilhelm2021 with respect to the cutoff radius of the truncated Coulomb metric and the auxiliary RI basis set size (here quantified by the $\Delta_{\mathrm{I}}$ metric threshold). As orbital basis set, we employ aug-SZV-MOLOPT-ae (top), aug-DZVP-MOLOPT-ae (middle) and aug-TZVP-MOLOPT-ae (bottom). The color map shows the mean absolute deviation of $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps for the same subset of 247 molecules from the GWGW5000 benchmark set used in Fig. 3, relative to a reference calculation using the aug-cc-pV5Z-RIFIT auxiliary RI basis set Haettig2005. Each row corresponds to an auxiliary RI basis set generated with a given $\Delta_{\mathrm{I}}$ metric threshold (from $10^{-2}$ to $10^{-5}$). Smaller errors are achieved with tighter RI thresholds and larger Coulomb cutoffs. A practical compromise is reached with a $\Delta_{\mathrm{I}}$ metric threshold of $10^{-4}$ and $r_{c}\hskip 1.69998pt{\geq}\hskip 1.69998pt7$ Å (error: 30 meV).
Fig. 5 summarizes the results. The color map shows the absolute deviation of the $G_{0}W_{0}$ HOMO-LUMO gaps (averaged over all 247 molecules) from the converged reference obtained with the large aug-cc-pV5Z-RIFIT auxiliary RI basis set Haettig2005 and cutoff $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt9$ Å. For the aug-SZV-MOLOPT basis set, at the loosest RI threshold ($10^{-2}$) and smallest cutoff ($r_{c}=3$ Å), the average error exceeds 300 meV. However, increasing the cutoff to $r_{c}=9$ Å reduces this error to $\sim$ 130 meV. For a larger auxiliary RI basis ($\Delta_{\mathrm{I}}$ threshold of $10^{-4}$), numerical accuracy is substantially improved: for $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt7$ Å, the deviation is 30 meV, and drops to 20 meV at $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt9$ Å. The best overall agreement with aug-cc-pV5Z-RIFIT is obtained for an even larger auxiliary RI basis ($\Delta_{\mathrm{I}}$ threshold of $10^{-5}$) with $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt9$ Å, where the average absolute error is reduced to 13 meV. The results for the aug-DZVP-MOLOPT and aug-TZVP-MOLOPT basis sets show better convergence properties than the aug-SZV-MOLOPT benchmark tests, which can be easily explained by the larger size of these basis sets. However, the overall convergence trends are very similar between all the basis sets. These results demonstrate that accurate low-scaling $GW$ calculations can be achieved using relatively compact auxiliary RI basis sets when paired with a sufficiently large Coulomb cutoff. For practical applications aiming at high numerical precision, we recommend a $\Delta_{\mathrm{I}}$ threshold of $10^{-4}$ and a cutoff radius of at least 7 Å, giving excellent balance between efficiency and accuracy ($\sim$ 30 meV).
Figure 6: GWGW calculations on nanographenes of increasing length (defined by the number of stripes $L\hskip 1.69998pt{=}\hskip 1.69998pt1,2,4,\dots,{\color[rgb]{0,0,0}512}$). (a) Nanographene geometry for $L\hskip 1.69998pt{=}\hskip 1.69998pt1,2,4,8$. Note that we put two hydrogen atoms at the center carbon atom at the zigzag edge to prevent for magnetic zigzag edge states. (b) Quasiparticle HOMO-LUMO gaps computed with $G_{0}W_{0}$ + Hedin’s shift @ PBE using different basis sets. Inset: deviation from the aug-DZVP-MOLOPT-ae basis set. (c) PBE LUMO eigenvalue, serving as a measure of LUMO diffuseness. (d) Condition number $\kappa(\mathbf{S})$ of the overlap matrix, computed from Eq. (11). (e) Computation time (in core hours) of the $G_{0}W_{0}$ calculations on Noctua2 (AMD Milan 7763) and Otus (AMD Turin 9655). The aug-MOLOPT basis sets exhibit low condition numbers and reduced computational cost, enabling stable and efficient calculations for nanographenes exceeding 9000 atoms. Details on the number of nodes used, wall time and memory consumption of the $GW$ calculations are listed in Table 3.
@@
To demonstrate the suitability of the generated augmented MOLOPT basis sets for large-scale applications, we perform $GW$ calculations on nanographenes of increasing size. Representative geometries are shown in Fig. 6a. For these systems, we employ the PBE functional Perdew1996 as the DFT starting point and Hedin’s shift (34) to avoid the higher cost of hybrid functionals during the SCF.
@@
We use auxiliary RI basis sets with a $\Delta_{\mathrm{I}}$ value below $1.5\cdot 10^{-2}$. The corresponding basis sizes are listed in Table 2. $G_{0}W_{0}$+H HOMO-LUMO gaps for the nanographenes are shown in Fig. 6b. For $L\hskip 1.69998pt{=}\hskip 1.69998pt1$ (9,10-Dihydroanthracene), basis set convergence is challenging: the minimal aug-SZV-MOLOPT-ae-mini basis underestimates the gap by approximately 1 eV. This is partly due to the small auxiliary RI basis used (cf. Fig. 5). For larger systems ($L\geq 8$), convergence improves significantly: all five basis sets agree within 50 meV (inset of Fig. 6b). This matches findings for two-dimensional materials Pasquier2025, where convergence within 100 meV was reached using the aug-SZV-MOLOPT basis.
We attribute the improved basis set convergence for larger structures to three factors: (i) larger systems offer more basis functions, increasing flexibility; (ii) the PBE LUMO energy decreases with $L$ (Fig. 6c), making the LUMO less diffuse in vacuum and thus easier to represent (as discussed for the decay length $\zeta_{n}\hskip 1.69998pt{\approx}\hskip 1.69998pt\hbar/{\sqrt{2m|\varepsilon_{n}|}}$ in Sec. 2.4). For a benchmark of the numerical precision of basis sets for $G_{0}W_{0}$ HOMO-LUMO gaps as function of the DFT LUMO energy, see the supporting information, Fig. S1 and S2 (and Fig. S3 for differences of excitation energies).; (iii) there may be cancellation of errors between an underconverged orbital and auxiliary RI basis set. Notably, error cancellation does not distort the size dependence: for $L\hskip 1.69998pt{=}\hskip 1.69998pt32$, 64, 128, all aug-SZV-MOLOPT (SR, mini) basis sets yield size-converged gaps consistently within 33 meV. This indicates robust $GW$ calculations for large systems. These results support two conclusions: (i) basis set convergence for nanostructures differs from that of small molecules and must be analyzed accordingly; (ii) further optimization of Gaussian basis sets for extended systems, in particular with pseudopotentials Goedecker1996, appears promising.
The condition number of the overlap matrix remains below $10^{7}$ for all augmented MOLOPT basis sets (Fig. 6d). In contrast, it exceeds $10^{13}$ for aug-cc-pVDZ. The computational cost is roughly reduced by a factor of 280 when using aug-SZV-MOLOPT-ae-mini instead of aug-cc-pVDZ (Fig. 6e). This aligns with expected $GW$ scaling of $N_{\text{bf}}^{2}N_{\text{RI}}^{2}$: According to Table 2, $N_{\text{bf}}$ and $N_{\text{RI}}$ decrease by factors of about 2.4 and 6.8, respectively, giving $2.4^{2}\hskip 1.19995pt{\cdot}\hskip 1.19995pt6.8^{2}\hskip 1.69998pt{\approx}\hskip 1.69998pt270$. Despite this enormous speedup, the gap difference between aug-SZV-MOLOPT-ae-mini and aug-cc-pVDZ is less than 10 meV for $L\hskip 1.69998pt{=}\hskip 1.69998pt16$. The small size of the aug-SZV-MOLOPT-ae-mini basis set enabled us to perform a $GW$ calculation on a nanographene with 9224 atoms, requiring only 34,300 core hours.